A referee tosses the coin to decide which team has to start the match. If the coin reaches a height of 1 m above the release point which was its initial speed?(assume for simplicity that g = 10 m/s2)
A: About 1 m/s
B: About 20 m/s
C: It is impossible to answer unless the time of flight is given
D: About 3.3 m/s
E: About 4.5 m/s
1
Expert's answer
2018-05-31T07:18:09-0400
V = √(V_0^2-2gh) , where V – velocity at the height of 1 m above the release point (0 m/s)
0 = V_0^2 – 2gh
V_0^2 = 2gh
V_0^2 = 2×10×1 = 20
V0 = 4.47 (m/s)
Answer
E: About 4.5 m/s